# InversionView: A General-Purpose Method for Reading Information from Neural Activations

### v2

@@ 5 Conclusion
We present InversionView, an effective method for decoding information from neural activations. By applying it in three case studies—character counting, IOI, and 3-digit addition—we showcase how it can reveal various types of information, thus facilitating reverse-engineering of algorithm implemented by neural networks. Moreover, we compare it with other interpretability methods and show its unique advantages. We also show that the results given by InversionView can in principle be interpreted automatically by LLMs, which opens up possibilities for a more automated workflow. This paper only explores a fraction of the opportunities this method offers. Future work could apply it to subspaces of residual stream (which can probably produce more insights), to larger models, or to different modalities such as vision.
@@ 6 Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102
@@ A.1 Observing Larger Neighborhoods is Important
We further prove this by causal intervention. We found that $x^{0,\mathrm{post}}_{+}$ has no effect on the model’s output. Concretely we patch $x^{0,\mathrm{post}}_{+}$ with its mean on the test set (mean ablation [39]) and for each prediction target (A1, A2 etc.), we compare 1) the KL divergence between the distribution before and after patching. 2) logit decrement ratio, which is the difference between the maximum logit value before patching and the logit value of the same target token after patching, divided by the former. E.g., 1.0 means the logit is reduced to zero (assuming it is originally positive). The results are shown in Table 1. We can see the effect of $x^{0,\mathrm{post}}_{+}$ is negligible.
@@
Logit decrement ratio $-2.3\times 10^{-5}$ $8.7\times 10^{-7}$ $1.4\times 10^{-5}$ $2.8\times 10^{-7}$
@@ A.2 Sampling with Decoder Model
In our experiments, we use both ways to control the generation, i.e., by adjusting the temperature and adding random noise to the query activation. We denote temperature as $\tau$ and noise coefficient as $\eta$. The noise vector consists of independent random variables sampled from the standard normal distribution and then multiplied by $\mbox{std}(\mathbf{z}^{q})\cdot\eta$ where $\mbox{std}(\cdot)$ stands for standard deviation. In our web application, we provide multiple sampling configurations: four configurations in which $\tau=\{0.5,1.0,2.0,4.0\}$ and $\eta=0.0$; one figuration named “Auto" which is sampled by following procedure: we iterate over a few predefined $\tau$ (ranging from 0.5 to 2.0) and $\eta$ (0.0 or 0.1) and sample a certain amount of inputs (e.g., 250) for each parameter combination. We then calculate the metric value for all inputs collected from different sampling configurations. We then randomly choose a small part of them (100) with different probability for in-$\epsilon$-preimage inputs and out-of-$\epsilon$-preimage inputs. We dynamically adjust the probability such that the in-$\epsilon$-preimage inputs account for 60%-80% of the chosen set of inputs (when this is possible).
@@ Appendix B Experimental Verification of Completeness
Appendix C Decoder Model Architecture
@@
The decoder is a decoder-only transformer. In order to condition the decoder on the query activation, the query activation is first passed through a stack of MLP layers to decode information depending on the activation sites and then made available to each attention layer of the decoder transformer.
Preprocessing Query Activation.
@@
The query activation $\mathbf{z}^{q}\in\mathbb{R}^{d}$ is first concatenated with a trainable activation site embedding $\mathbf{e}_{act}\in\mathbb{R}^{d_{site}}$, producing the intermediate representation $\mathbf{z}^{(0)}=[\mathbf{z}^{q};\mathbf{e}_{act}]$. We chose $d_{site}$ to be the number of possible activation sites in the training set. The result $\mathbf{z}^{(0)}$ is then fed through 6 MLP layers with residual connections:
$\mathbf{z}^{(i+1)}=\mbox{MLP}(\mbox{LN}(\mathbf{z}^{(i)}))+\mathbf{z}^{(i)}$ (2)
@@
where LN represents layer normalization. For each MLP layer, the input, hidden and output dimensions are $d+d_{site}$, $d$, and $d+d_{site}$, respectively. The activation function is ReLU. There is also a final layer normalization, $\mathbf{z}^{(fn)}=\mbox{LN}(\mathbf{z}^{(6)})$.
Providing Query Activation to Decoder.
@@
As we want to make the query activation available to each attention layer of the decoder, we separately customize it to the needs of each layer using a linear layer. That is, for each layer of the decoder transformer (indexed by $\ell$), we define a linear layer $\mbox{L}^{(\ell)}:\mathbb{R}^{d+d_{site}}\rightarrow\mathbb{R}^{d_{decoder}}$ and a layer normalization $\mbox{LN}^{(\ell)}$, where $d_{decoder}$ is the model dimension of the decoder model:
$\mathbf{z}^{(l)}=\mbox{LN}^{(\ell)}(\mbox{L}^{(\ell)}(\mathbf{z}^{(fn)}))$ (3)
@@
In the $\ell$-th layer ($\ell=1,2,\dots$) of the decoder, $\mathbf{z}^{(\ell)}$ is concatenated with the input of the attention layer along the length axis before computing keys and values, so that each attention head can also attend to $\mathbf{z}^{(\ell)}$ in addition to the context tokens. This means that each head in the $\ell$-th layer of the decoder, instead of attending to $x^{(\ell)}_{1},x^{(\ell)}_{2},...,$ now computes its attention distribution over $\mathbf{z}^{(l)},x^{(\ell)}_{1},x^{(\ell)}_{2},...$. Here $x^{(\ell)}_{t}$ is the residual stream corresponding to the $t$-th token input into the $\ell$-th layer. We note that there are other possible choices for conditioning the generation on the activation, and we didn’t optimize this choice thoroughly.
@@
The decoder transformer has 2 layers, 4 heads per layer, and a model dimension of 256. The attention dropout rate is 0. Other settings are the same as the GPT-2. We use the same architecture for all 3 tasks in the paper.
@@ D.1 Implementation Details
With regard to the decoder model, the architecture is described in C. We select $x^{0,\mathrm{pre}},x^{i,\mathrm{mid}},x^{i,\mathrm{post}},a^{i,j},m^{i}$ as the set of activation sites we are interested in, where $i\in\{0,1\},j\in\{0\}$, and $m$ denotes MLP layer output. The query activation is sampled from those activations corresponding to only the target character and colon. By default, we sample query activation with equal probability of activation sites, as is done when training the decoder for the other two case studies. During training, we test the generation quality quantified by the ratio of samples (temperature=1) that are inside the $\epsilon$-preimage ($\epsilon=0.1$). For those activation sites for which the decoder has a low generation quality, we increase their probability of being sampled in the training data.
@@ D.2 More Examples of InversionView
See Figures 12 and 13. Here, we show results similar to Figure 2 for other query inputs.
@@ E.1 Implementation Details
The architecture of the decoder model is the same as before, as described in Appendix C. The set of activation sites the decoder is trained on consists of the output of all attention heads and MLP layers (no residual stream). Note that when producing query activation using GPT-2, we always add the “<|<|< |endoftext|>|>| >” token as the BOS token. We do so as during the training of GPT-2 this or multiple such tokens that usually appear in the previous context can be used as a BOS token, which is possibly important to the model’s functioning. We use a new token “[EOS]” as the EOS token when training the decoder. The query activation is sampled uniformly from all positions excluding EOS and padding tokens, and uniformly from all activation sites the decoder is trained for.
@@ E.2 More Examples of InversionView
See Figures 17, 18, 19.
Figure 17: $\epsilon$-preimage of S-Inhibition Head 7.3. The relative position of S1 – but not its identity, is contained in the head output. That means, in the samples within the $\epsilon$-preimage, S1 always appears before the IO. While the relative position is encoded, the absolute position can vary, as can the identities of the names.
@@ E.4 Choice of Distance Metric in IOI
In Figure 21 and 22 we show two examples of using normalized Euclidean distance as the distance metric. Readers can see more examples on our web application. We see that the samples generated by the decoder tend to have larger distances under normalized Euclidean distance. Nonetheless, we can still see the top-ranked samples show meaningful commonalities – indeed if we set $\epsilon=0.4$, we obtain the same interpretation as the one we obtain based on Figure 3(b) and 17.
@@ F.1 Implementation Details
With regard to the decoder model, we select $x^{0,\mathrm{pre}},x^{i,\mathrm{mid}},x^{i,\mathrm{post}}a^{i,j}$ as the set of activation sites we are interested, where $i\in\{0,1\},j\in\{0,1,2,3\}$. When sampling training data, we select an activation site and a token position uniformly at random.
@@ F.2 More Examples of InversionView
See Figures 24 to 27.333We have also tried value-weighted attention pattern [23]; it makes a negligible difference, so we always show the original attention pattern in our paper.
@@ F.5 Causal Intervention: Details and Full Results
Activation patching results for other cases are shown in Figures 30 to 34.
@@ Appendix G Decoder Likelihood Difference
Appendix H Notes on Attention and Path Patching
H.1 InversionView Reveals Information from Un-attended Tokens

### v3

@@ 6 Conclusion
We present InversionView, an effective method for decoding information from neural activations. By applying it in four case studies—character counting, IOI, 3-digit addition, and factual recall—we showcase how it can reveal various types of information, thus facilitating reverse-engineering of algorithm implemented by neural networks. Moreover, we compare it with other interpretability methods and show its unique advantages. We also show that the results given by InversionView can in principle be interpreted automatically by LLMs, which opens up possibilities for a more automated workflow. This paper only explores a fraction of the opportunities this method offers. Future work could apply it to subspaces of residual stream (which can probably produce more insights), to larger models, or to different modalities such as vision.
@@ 7 Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102. We thank anonymous reviewers for their encouraging and constructive feedback.
@@ A.1 Observing Larger Neighborhoods is Important
We further prove this by causal intervention. We found that $x^{0,\mathrm{post}}_{+}$ has no effect on the model’s output. Concretely we patch $x^{0,\mathrm{post}}_{+}$ with its mean on the test set (mean ablation [50]) and for each prediction target (A1, A2 etc.), we compare 1) the KL divergence between the distribution before and after patching. 2) logit decrement rate, which is the difference between the maximum logit value before patching and the logit value of the same target token after patching, divided by the former. E.g., 1.0 means the logit is reduced to zero (assuming it is originally positive). The results are shown in Table 1. We can see the effect of $x^{0,\mathrm{post}}_{+}$ is negligible.
@@
Logit decrement rate $-2.3\times 10^{-5}$ $8.7\times 10^{-7}$ $1.4\times 10^{-5}$ $2.8\times 10^{-7}$
@@ A.2 Sampling with Decoder Model
In our experiments, we use both ways to control the generation, i.e., by adjusting the temperature and adding random noise to the query activation. We denote temperature as $\tau$ and noise coefficient as $\eta$. The noise vector consists of independent random variables sampled from the standard normal distribution and then multiplied by $\mbox{std}(\mathbf{z}^{q})\cdot\eta$ where $\mbox{std}(\cdot)$ stands for standard deviation. In our web application, we provide multiple sampling configurations: four configurations in which $\tau=\{0.5,1.0,2.0,4.0\}$ and $\eta=0.0$ (only for addition task); one figuration named “Auto" which is sampled by following procedure: we iterate over a few predefined $\tau$ (ranging from 0.5 to 2.0) and $\eta$ (0.0 or 0.1) and sample a certain amount of inputs (e.g., 250) for each parameter combination. We then calculate the metric value for all inputs collected from different sampling configurations. We then randomly choose a small part of them (100) with different probability for in-$\epsilon$-preimage inputs and out-of-$\epsilon$-preimage inputs. We dynamically adjust the probability such that the in-$\epsilon$-preimage inputs account for 60%-80% of the chosen set of inputs (when this is possible). Note that we use different noise in factual recall task, which will be described later in Appendix H.4.
@@ Appendix B Experimental Verification of Completeness
Appendix C Decoder Model
@@
The decoder model is basically a decoder-only transformer combined with some additional MLP layers. In order to condition the decoder on the query activation, the query activation is first passed through a stack of MLP layers to decode information depending on the activation sites and then made available to each attention layer of the transformer part of the decoder, as depicted in Figure 11.
Figure 11: The decoder model architecture used in this paper. The query activation is processed by a stack of MLP layers before being available as part of the context in attention layers. We use transparent blocks to represent model components inherited from original decoder-only transformer model.
Processing Query Activation.
@@
The query activation $\mathbf{z}^{q}\in\mathbb{R}^{d}$ is first concatenated with a trainable activation site embedding $\mathbf{e}_{act}\in\mathbb{R}^{d_{site}}$, producing the intermediate representation $\mathbf{z}^{(0)}=[\mathbf{z}^{q};\mathbf{e}_{act}]$. We chose $d_{site}$ to be the number of possible activation sites in the training set. The result $\mathbf{z}^{(0)}$ is then fed through multiple MLP layers (each layer indexed by $p\in\{0,1,\cdots,P-1\}$) with residual connections:
$\mathbf{z}^{(p+1)}=\mbox{MLP}(\mbox{LN}(\mathbf{z}^{(p)}))+\mathbf{z}^{(p)}$ (2)
@@
where LN represents layer normalization. For each MLP layer, the input, hidden and output dimensions are $d+d_{site}$, $d$, and $d+d_{site}$, respectively. The activation function is ReLU. There is also a final layer normalization, $\mathbf{z}^{(fn)}=\mbox{LN}(\mathbf{z}^{(P)})$.
Integrating Query Activation.
@@
As we want to make the query activation available to each attention layer of the decoder, we separately customize it to the needs of each layer using a linear layer. That is, for each layer of the transformer part of the decoder (indexed by $\ell$, so $\ell\in\{0,1,\cdots,L-1\}$), we define a linear layer $\mbox{Linear}^{(\ell)}:\mathbb{R}^{d+d_{site}}\rightarrow\mathbb{R}^{d_{decoder}}$ and a layer normalization $\mbox{LN}^{(\ell)}$, where $d_{decoder}$ is the model dimension of the decoder model:
$\hat{\mathbf{z}}^{(\ell)}=\mbox{LN}^{(\ell)}(\mbox{Linear}^{(\ell)}(\mathbf{z}^{(fn)}))$ (3)
@@
We add superscript $(\ell)$ to the model components to emphasize they are layer-specific. In the $\ell$-th layer of the transformer part, $\hat{\mathbf{z}}^{(\ell)}$ is concatenated with the input of the attention layer along the length axis before computing keys and values, so that each attention head can also attend to $\hat{\mathbf{z}}^{(\ell)}$ in addition to the context tokens. This means that each head in the $\ell$-th layer, instead of attending to $x^{(\ell)}_{1},x^{(\ell)}_{2},...,$ now computes its attention weights over $\hat{\mathbf{z}}^{(l)},x^{(\ell)}_{1},x^{(\ell)}_{2},...$. Here $x^{(\ell)}_{t}$ is the residual stream corresponding to the $t$-th token input into the $\ell$-th layer.
Motivation for Architecture Design.
At the beginning, we train separate decoders for each activation site, but this is not very scalable when there are many activation sites. In the architecture above mentioned, we use activation site embedding $\mathbf{e}_{act}$ as a signal to trigger different processing, and functions or model components that are needed for all activation sites are shared. Similar reason applies to the linear layer $\mbox{Linear}^{(\ell)}$, we expect the $\mathbf{z}^{q}$ should be transformed differently for each transformer layer, but having separate MLP stacks to process $\mathbf{z}^{q}$ for each layer would largely increase the number of parameters. In preliminary experiments, we also try “encoder-decoder” attention layer. That is, instead of providing query activation in self attention layer, we add new attention layers that analogous to the encoder-decoder attention layer in original transformer architecture, where each token can attend to the processed query activation as well as a blank representation (similar to the function of “BOS", so that “no-op" is possible). However, we do not find significant difference between this design and the aforementioned one. Therefore, other than adding components for processing $\mathbf{z}^{q}$, we do not modify the decoder-only transformer architecture, so that we can also choose to use pretrained models. We note that there are other possible choices for conditioning the generation on the activation, and we didn’t optimize this choice thoroughly.
@@
Regarding processing query activation, the decoder has 6 MLP layers, i.e., $P=6$. The decoder model has 2 transformer layers ($L=2$), 4 heads per layer, and a model dimension of 256. The attention dropout rate is 0. Other settings are the same as the GPT-2. We use the this architecture for the 3 tasks—character counting, IOI, and 3-digit addition— in the paper. Regarding the factual recall task, we use the same architecture for processing query activation, i.e. $P=6$, and use pretrained GPT-2 Medium (24 layers) as the transformer part of the decoder.
Training Details.
We construct the training dataset by feeding in-domain inputs to the probed model, and collect activations from random activation sites and random position as query activations (the choice of activation sites and position is specific to each task and is described later), we also record the activation site they come from. For each input, we could obtain many possible training examples because of many choices of activation sites and position. So we do not iterate over all possible training examples. We sample certain amount of examples to train the decoder for 1 epoch, using constant learning rate of 0.0001 and AdamW optimizer with weight decay of 0.01. Other details (e.g., amount of examples, training steps) are task-specific, and can be found later in their own section.
During training, we regularly calculate the in-preimage rate, which serves as a proxy for generation quality. Concretely, for a fixed set of query activations used for testing, the decoder generates samples with temperature=1, we then compute the fraction of samples inside of $\epsilon$-preimage (with $\epsilon=0.1$, $D$ as normalized Euclidean distance). The rate is calculated for each activation site. We usually observe difference between the ratios for each activation site, indicating some inverse mappings are easier to learn (we also observe these activation sites tend to have clearer information). Overall, we usually see a continuous improvement on the average rate during training.
@@ D.1 Implementation Details
With regard to the decoder model, the architecture is described in C. We select $x^{0,\mathrm{pre}},x^{i,\mathrm{mid}},x^{i,\mathrm{post}},a^{i,j},m^{i}$ as the set of activation sites we are interested in, where $i\in\{0,1\},j\in\{0\}$, and $m$ denotes MLP layer output. The query activation is sampled from those activations corresponding to only the target character and colon. We sample 100 million training examples (all activation sites are included) and train the decoder with batch size of 512, resulting in roughly 200K steps. During training, as we mentioned before in C, we test the generation quality by measuring in-preimage rate. For those activation sites for which the decoder has a low generation quality, we increase their probability of being sampled in the training data. The final in-preimage rate averaged across activation sites is 67.7%.
@@ D.2 More Examples of InversionView
See Figures 13 and 14. Here, we show results similar to Figure 2 for other query inputs.
@@ E.1 Implementation Details
The architecture of the decoder model is the same as before, as described in Appendix C. The set of activation sites the decoder is trained on consists of the output of all attention heads and MLP layers (no residual stream). Note that when producing query activation using GPT-2, we always add the “<|<|< |endoftext|>|>| >” token as the BOS token. We do so as during the training of GPT-2 this or multiple such tokens that usually appear in the previous context can be used as a BOS token, which is possibly important to the model’s functioning. We use a new token “[EOS]” as the EOS token when training the decoder. The query activation is sampled uniformly from all positions excluding EOS and padding tokens, and uniformly from all activation sites the decoder is trained for. We sample 20 million training examples (all activation sites are included) and train the decoder with batch size of 256, resulting in roughly 80K steps. The final average in-preimage rate is 58.0%, despite that the decoder is trained for 157 activation sites.
@@ E.2 More Examples of InversionView
See Figures 18, 19, 20.
Figure 18: $\epsilon$-preimage of S-Inhibition Head 7.3. The relative position of S1–but not its identity–is contained in the head output (together with some template information). That means, in the samples within the $\epsilon$-preimage, S1 always appears before the IO. While the relative position is encoded, the absolute position can vary, as can the identities of the names.
@@ E.4 Choice of Distance Metric in IOI
In Figure 22 and 23 we show two examples of using normalized Euclidean distance as the distance metric. Readers can see more examples on our web application. We see that the samples generated by the decoder tend to have larger distances under normalized Euclidean distance. Nonetheless, we can still see the top-ranked samples show meaningful commonalities – indeed if we set $\epsilon=0.4$, we obtain the same interpretation as the one we obtain based on Figure 3(b) and 18.
@@ F.1 Implementation Details
With regard to the decoder model, we select $x^{0,\mathrm{pre}},x^{i,\mathrm{mid}},x^{i,\mathrm{post}}a^{i,j}$ as the set of activation sites we are interested, where $i\in\{0,1\},j\in\{0,1,2,3\}$. When sampling training data, we select an activation site and a token position uniformly at random. We sample 100 million training examples (all activation sites are included) and train the decoder with batch size of 512, resulting in roughly 200K steps. The final average in-preimage rate is 88.6%.
@@ F.2 More Examples of InversionView
See Figures 25 to 28.555We have also tried value-weighted attention pattern [31]; it makes a negligible difference, so we always show the original attention pattern in our paper.
@@ F.5 Causal Intervention: Details and Full Results
Activation patching results for other cases are shown in Figures 31 to 35.
@@ Appendix G Decoder Likelihood Difference
Appendix H Factual Recall: Detailed Findings
H.1 Background
We use InversionView to investigate how GPT-2 XL (1.5B parameters) performs the task of recalling factual associations. In recent research [24, 11, 20], the task has been formalized in terms of triples $(s,r,a)$, where $s$, $r$, $a$ are subject, relation, attribute respectively. The model should predict $a$ based on input containing $s$ and $r$ in natural language. E.g., given the prompt: “LGA 775 is created by", the model is expected to predict “Intel". In this section of the paper, we refer the last token as END, and the last subject token as SUBJ. In the above example, END is “by" and SUBJ is “75". Geva et al. [24] found a high-level pattern that GPT-2 XL uses in solving this task: Early-middle MLP layers at SUBJ integrate information about the subject into its residual stream. Meanwhile, the relation information is incorporated into END’s residual stream through early attention layers. In upper layers, the correct attribute is moved from SUBJ to END’s residual stream by attention layers.
H.2 Selecting Activation Sites
As we mentioned earlier, our goal here is not to provide a full interpretation of the computations performed to solve this task; rather, it is to test whether InversionView can successfully produce interpretable results on larger models. As the factual recall task is complex and involves many model components [55, 11], we decide to focus on 25 attention heads in the upper part of the model (layer 24-47) that contribute most frequently to the final prediction. We select these activation sites because the attribute retrieval tends to happen there [24] and we expect more abstract information in $\epsilon$-preimage. Here, we note that preliminary experiments revealed that it is necessary to restrict the number of activation sites that the decoder is trained for, given a limited compute budget, as more activation sites require the decoder to learn more a complex overall inverse mapping from activations to inputs. Scaling the approach to apply to many activation sites simultaneously is left to future work.
Concretely, we use the attribution method introduced in [18] to estimate the head importance at END position. Formally, given $y=z_{1}+\cdots z_{m}$, the importance of each term $z_{j}$ to the sum $y$ is estimated as follows:
$\text{importance}(z_{j},y)=\frac{\text{proximity}(z_{j},y)}{\sum_{k}\text{proximity}(z_{k},y)},$ (4)
$\text{proximity}(z_{j},y)=\max(-\|z_{j}-y\|_{1}+\|y\|_{1},0).$ (5)
[…]
There are multiple methods that can be used to find important attention heads [37, 38, 46, 28], because we do not have strict requirements for finding the most important heads, we choose this one because of its simplicity.
H.3 Decoder Training
To train the decoder model, we collect text from 3 datasets, including the factual statements from CounterFact [36]777 We make use of all prompts of each data point. Besides the ”prompt”, we also include ”paraphrase_prompts”, ”neighborhood_prompts”, and ”attribute_prompts” concatenated with their corresponding answer. and BEAR [52]  [^8] We use $\text{BEAR}_{\text{big}}$, as well as general text from MiniPile [29]. The factual statements we used are complete sentences containing the final attribute. Note that in CounterFact, for each attribute, there are always multiple different subjects associated with it. So if the information contained is solely the attribute, a well-trained decoder should generate different subjects. Regarding MiniPile, we randomly select 10% of it, and split the text into sentences [9] and remove sentences longer than 100 characters.
Importantly, for each head we selected, we extract a separate subset of sentences from the collected data on which the head is “activated". By being “activated" we mean the attention weight on BOS token is less than 0.6999we do not experiment with other values. We find this is important based on preliminary experiments, because in many cases attention heads in higher layers execute “no-op" by exclusively attending to BOS, resulting in attention output containing no information. Training decoder on these activations discourages it to learn to read information from the query activation. So each head correspond to a subset of text, on which the head output will be captured to create training data. The number of sentences in subsets ranges from 0.6 to 2.6 million.
The training data for decoder consists of two parts, each of which accounts for 50%. In one part, the activations are taken from GPT-2 XL when processing factual statements from CounterFact and BEAR, and activations correspond to the END position (the sentence structure is provided in these datasets). In the other part, the activations correspond to text from all 3 datasets (thus mainly composed of MiniPile text), and correspond to the position with least attention weight on BOS token. By doing so, we emphasize the importance of factual statements domain while covering a large variety of text.
We use GPT-2 Medium as the backbone model for decoder. Concretely, the newly added components in the decoder architecture (e.g. those parts responsible for processing query activation as shown in Figure 11) are trained from scratch, but the transformer layers in the decoder are equipped with pretrained weights. In this way, we also make use of the existing capacity of language models. Note that in other case studies, we use a small 2 layer transformer as decoder. The reason why we use a much larger decoder for this task is we expect a much more complex inverse mapping to be learned. Specifically, we expect $\epsilon$-preimage for some activation sites to contain multiple different subjects sharing a certain attribute, the decoder needs to generate these subjects given the attribute. So it needs enough capacity to memorize the knowledge.
Similar as before, we always add the “<|<|< |endoftext|>|>| >” token as the BOS token when capturing query activation from GPT-2 XL. We again use a new token “[EOS]” as the EOS token when training the decoder. We sample query activation with equal probability of activation sites. We sample 32 million training examples (all activation sites are included) and train the decoder with batch size of 512, resulting in roughly 60K steps. The final average in-preimage rate is 36.0%.
H.4 Sampling from Decoder
Similar as before, when using decoder to generate samples, we perturb the query activation before feeding it to the decoder in order to cover the neighbourhood around the query. In this larger model with high-dimensional activations, we found it useful to craft this process more carefully and use different disturbance. In preliminary experiments, we found that cosine similarity usually produces more interpretable $\epsilon$-preimages, therefore, we randomly sample activations that have a certain angle $\theta$ to the query activations, and then randomly scale it so that its magnitude ranges from $e^{-1}\cdot\|\mathbf{z}^{q}\|_{2}$ to $e\cdot\|\mathbf{z}^{q}\|_{2}$. We repeat this process for different $\theta$ values with in the range $\cos\theta\in[0.75,0.9]$. To further encourage diversity, we lower the probability of tokens that have already appeared too many times in the generated samples. [^10] We chose a simple and heuristic approach, without any attempt at optimizing it: Any token that has appeared $>5$ times in the same position, while appearing less than 50K times in the miniPile training set, has its probability set to zero at this position.
We use the first 200 examples of the aforementioned subset of CounterFact (the subset containing examples known by GPT-2 XL) as query inputs and generate samples. The results are also available in our web application.
H.5 Observation
Choice of Threshold.
We use the same distance metric as in IOI task, and we set the threshold $\epsilon=0.25$. As we mentioned earlier, larger threshold produces more coarse-grained information. Because we observe that the trained decoder in many cases does not generate enough and diverse samples within a close distance (e.g., 0.1), we increase the threshold in order to draw more reliable conclusions.
A priori, the sparsity of sampled $\epsilon$-preimages at smaller thresholds may reflect multiple possibilities. One is that there indeed are no close neighbors for the query input. For example, if the subject’s text is copied, then only samples containing the same subject or the same token will lie in $\epsilon$-preimage. Such a phenomenon may generally reflect the high-dimensional geometry of larger models. Another one is that the decoder is not complete; and that it would either require more training or more capacity, or more training data. [^11] For example, if the query activation contains information “related to Amazon and related to audio”, there are only a few satisfactory subjects in the dataset (“Audible.com”, “Amazon Music”). Training larger decoders on more data would mitigate this problem, which can be done in future work. Nevertheless, in some cases, we do observe a substantial number of diverse samples even within $\epsilon=0.1$, allowing us to infer information at a higher level of granularity in these cases.
Figure 37: $\epsilon$-preimage for head output $a^{24,24}$, showing encoded information is the relation. On the left, the relation “developed by" appears throughout the $\epsilon$-preimage. On the right, the relation “domain of work is" is consistent $\epsilon$-preimage.
Some heads have fixed behavior
[…]
Figure 38: $\epsilon$-preimage for head output $a^{29,9}$, which contains information about the subject. On the left, Czech-related words are common throughout the $\epsilon$-preimage, e.g., “Czech", “Prague", “Bohem…", so the information contained is the subject is “Czech-related". On the right, “Nokia" is shared in $\epsilon$-preimage, so the information is simply “Nokia".
[…]
Figure 39: $\epsilon$-preimage for head output $a^{35,19}$, which exhibit different information on different inputs. On the left, Philippines-related words are common throughout the $\epsilon$-preimage. For example, “Cebu": a city in Philippines, “TV Patrol Southern Tagalog": a TV program in Philipines, “Enrile": A municipality of Philippines and a surname of many famous Filipinos, “GMA": a Philippine television channel / broadcasting company. So the information contained is the subject is “Philippines-related". Note that because the samples are generated, the fact stated in sample is not true in many cases. On the right, the relation “performs on" is encoded.
Other heads exhibit a mix of behaviors
The other heads among the 25 heads we inspect, when “activated”, move information about subject or relation (in some cases, both). Which behavior is exhibited varies between inputs. Figure 39 shows one of these heads. On the left of the figure, the head moves information about subject, while on the right it moves information about the relation.
Figure 40: Examples showing relation-agnostic retrieval. On the left, the information encoded is “soccer", which is indeed the requested attribute. However, the first sample shows this is not dependent on the relation, since the “soccer" is still retrieved when relation is “speaks language". On the right, the information “audio-related" is encoded, while the relation in the query input is “owned by".
Relation-agnostic retrieval
In the factual recall task, only one specific attribute is sought. Geva et al. [24] found evidence that the model “queries” the residual stream at SUBJ for the specific attribute asked by the relation part of the prompt, and the representation at END can be viewed as such a relation query. In other words, the attribute extracted from the subject representation depends on the relation. However, we do not observe reliable evidence for this phenomenon among the 25 heads we inspect. In the figures we have shown so far, we can see the commonality shared between samples in $\epsilon$-preimage is not the attribute requested in the prompt. This still holds if one reduces the threshold. Figure 40 shows two more examples showing this characteristic. In general, we find that the information moved by the 25 attention heads tends to be the most important attribute or the “main” attribute about the subject. On the left of Figure 40, the most important attribute coincides with the requested attribute. We know this is a coincidence because the closest sample in $\epsilon$-preimage has a different relation. Note that other attributes about the subject could be moved by heads that we have not studied.
Figure 41: Examples showing different attributes of the same subject are extracted by different heads. In the query input, “Joseph Schumpeter" is an Austrian political economist. On the left, the information encoded is “economist". On the right, the information is about language/nationality (areas around Austria). Again, we emphasize that the facts stated in the sample are not necessarily true.
In addition, while heads that attend to subject usually move the most important attribute, we do observe sometimes different attributes are moved by different heads. In Figure 41, we show that information about profession and information about language/nationality of the subject are extracted by two heads. In fact, we observe that head $h^{31,0}$ tends to extract language/nationality information in general. But we do not find other obvious pattern of attribute category extracted by other heads.
Our findings echo those from [11], who argue that the primary mechanism for the factual recall task is additive. With our running example “LGA 775 is created by”, simply speaking, some heads promote attributes associated with the subject (chip, hardware, Intel, etc.), some heads promote attributes associated with the relation (Apple, Nintendo, Intel, etc). When the results from these independent mechanisms are added together, the intersection (Intel) will stand out. Therefore, the model can solve the task by adding two simple circuits, while humans find Q-composition [17] (i.e., relation information is used as queries in attention heads) more intuitive. From another perspective, this mechanism implies vector arithmetic. Instead of vector addition in vocabulary space, we can think of it as first summing two vectors (e.g., the output of subject heads and the output of relation heads) and then projecting them to vocabulary space.
The evidence found by Geva et al. [24] can also be explained by this mechanism. They use DLA to inspect attention layers’ updates to END’s residual stream, and find that the token most strongly promoted by each update usually matches the attribute predicted at the final layer. In other words, after projecting attention layer’s output to vocabulary space, the top-1 token is usually the exact requested attribute. Because their experiments study attention layer’s output as a whole, instead of individual heads, an alternative explanation is the additive mechanism mentioned above. Importantly, because we only inspect 25 heads, other mechanisms including Q-composition are also possible. Moreover, different mechanism can exist in other models, readers should not draw strong conclusions.
H.6 More Examples of InversionView
See Figure 42, Figure 43. Interestingly, when inspecting the $\epsilon$-preimage shown on the right of Figure 43, we also find an example showing InversionView can detect a flaw of the model. Our interpretation for this activation is “Canada-related", and we can see “York University” (which is in Canada) inside the preimage, while “University of York" (a different university, located in England) is outside of the preimage. However, we find that the “U of York" is inside the preimage.
Checking the model’s prediction about “U of York", we find that the model indeed believes “U of York" is in Canada. More specifically, given the prompt “U of York is located in", the top-12 predictions for the next token are: the, Toronto, York, downtown, a, one, New, central, Canada, North, Scarborough, London. On the contrary, with the prompt “University of York is located in" the model’s top-12 predictions are: the, York, North, East, central, Yorkshire, north, a, northeast, London, England, northwest.
Figure 42: $\epsilon$-preimage showing information about the subject moved by the attention head. On the left, the information is “cpu/computer-hardware-related". On the right, the information is “island country". Note that some statements are not correct.
Figure 43: $\epsilon$-preimage showing different information of subject is moved by different head. On the left, the information is superficial text content “York" and/or “University". Samples containing these words are in $\epsilon$-preimage, such as “University of York", even though it is a different university located in England. This can be confirmed by the fact that “Wellington, Ontario" is far away. On the right, the information is “Canada-related", which is more high-level. We can also see “University of York" is outside of the preimage.
Appendix I Notes on Attention, Path Patching and DLA
I.1 InversionView Reveals Information from Un-attended Tokens